Given the equation: A = \frac{B}{C}

solve for c Solve for C, or rather, what is C equal to (in terms of A and B

solve for b Solve for B, or rather, What is B equal to
Answer #1

A = B/C

C = B/A

And B = A*C

Since A = B divided by C, C will be equal to B divided by A

and hence, B equals A multiplied by C